# Cliqueformer: Model-Based Optimization with Structured Transformers

### v1

@@ Abstract
Expressive large-scale neural networks enable training powerful models for prediction tasks. However, in many engineering and science domains, such models are intended to be used not just for prediction, but for design—e.g., creating new proteins that serve as effective therapeutics, or creating new materials or chemicals that maximize a downstream performance measure. Thus, researchers have recently grown an interest in building deep learning methods that solve offline model-based optimization (MBO) problems, in which design candidates are optimized with respect to surrogate models learned from offline data. However, straightforward application of predictive models that are effective at predicting in-distribution properties of a design are not necessarily the best suited for use in creating new designs. Thus, the most successful algorithms that tackle MBO draw inspiration from reinforcement learning and generative modeling to meet the in-distribution constraints. Meanwhile, recent theoretical works have observed that exploiting the structure of the target black-box function is an effective strategy for solving MBO from offline data. Unfortunately, discovering such structure remains an open problem. In this paper, following first principles, we develop a model that learns the structure of an MBO task and empirically leads to improved designs. To this end, we introduce Cliqueformer—a scalable transformer-based architecture that learns the black-box function’s structure in the form of its functional graphical model (FGM), thus bypassing the problem of distribution shift, previously tackled by conservative approaches. We evaluate Cliqueformer on various tasks, ranging from high-dimensional black-box functions from MBO literature to real-world tasks of chemical and genetic design, consistently demonstrating its state-of-the-art performance.
@@ 1 Introduction
Most of the common use cases of deep learning (DL) so far have taken the form of prediction tasks (Hochreiter & Schmidhuber, 1997; He et al., 2016; Krizhevsky et al., 2017; Vaswani et al., 2017). However, in many applications, e.g. protein synthesis or chip design, we might want to use powerful models to instead solve optimization problems. Clearly, accurate predictions of a target score of an object could be used to find a design of the object that maximizes that score. Such a methodology is particularly useful in engineering problems in which evaluating solution candidates comes with big risk. For example, synthesizing a proposed protein requires a series of wet lab experiments and induces extra cost and human effort (Gómez-Bombarelli et al., 2018; Brookes et al., 2019). Thus, to enable proposing de-novo generation of strong solution candidates, researchers have drawn their attention to offline black-box optimization (BBO), often referred to as model-based optimization (MBO). In this paradigm, first, a surrogate model of the score is learned from offline data. Next, a collection of designs is trained to maximize the surrogate, and then proposed as candidates for maximizers of the target score (Gómez-Bombarelli et al., 2018; Kumar et al., 2021).
@@
Figure 1: The first building block of the Lat. RBF tasks are 3D radial-basis functions (left). These functions are applied to triplets arranged in a chain of triangles FGM (center) and linearly mixed. Then, an observable designs are produced with non-linear transformations of the chain and, together with their values, form a dataset. We show the score (right) of our structure-learning Cliqueformer and structure-oblivious COMs (Trabucco et al., 2021), against the dimension of Lat. RBF functions, modulated only by varying the number of triangles. Cliqueformer, unlike COMs, sustains strong performance across all dimensions. More results in Section 5.
In contrast to previous works, in our paper, we develop a scalable model that tackles MBO by learning the structure of the black-box function through the formalism of functional graphical models. Our architecture aims to solve MBO by 1) decomposing the predictions over the cliques of the function’s FGM, and 2) enforcing the cliques’ marginals to have wide coverage with our novel form of the variational bottleneck (Kingma & Welling, 2013; Alemi et al., 2016). However, building upon our Theorem 2, we do not follow grudzien2024functional during the FGM discovery step, and instead subsume it in the learning algorithm. To enable scaling to high-dimensional problems and large datasets, we employ the transformer backbone (vaswani2017attention). Empirically, we demonstrate that our model, Cliqueformer, inherits the scalability guarantees of MBO with FGM (see Figure 1). We further demonstrate its superiority to baselines in a suite of tasks with latent radial-basis functions (grudzien2024functional) and real-world chemical (hamidieh2018data) and DNA design tasks (trabucco2022design; uehara2024understanding).
@@ 2 Preliminaries
In this section, we provide the necessary background on model-based optimization. Additionally, we cover the basics of functional graphical models on top of which we build Cliqueformer.
2.1 Model-based Optimization
We consider a model-based optimization problem, where we are given a dataset $\mathcal{D}=\{{\mathbf{x}}^{i},{\textnormal{y}}^{i}\}_{i=1}^{N}$ of examples ${\mathbf{x}}\in\mathcal{X}$, following distribution $p({\mathbf{x}})$, and their values ${\textnormal{y}}=f({\mathbf{x}})\in\mathbb{R}$ under an unkown (black-box) function $f:\mathcal{X}\rightarrow\mathbb{R}$. Our goal is to optimize this function offline—to find its maximizer
@@ 2.1 Model-based Optimization
by only using information provided in $\mathcal{D}$ (kumar2020model). Sometimes, a more general objective in terms of a policy over $\pi({\mathbf{x}})$ is also used, $\eta({\mathbf{x}})=\mathbb{E}_{{\mathbf{x}}\sim\pi}[f({\mathbf{x}})]$. In either form, unlike in Bayesian optimization, we cannot make additional queries to the black-box function (brochu2010tutorial; kumar2020model). This formulation represents settings in which obtaining such queries is prohibitively costly, such as tests of new chemical molecules or of new hardware architectures (kim2016pubchem; kumar2021data).
@@
where $\text{Reg}(\theta)$ is an optional regularizer. Classical methods choose $\text{Reg}(\theta)$ to be identically zero, while conservative methos use the regularizer to bring the values of examples out of $\mathcal{D}$ down. For example, the regularizer of Conservative Objective Model’s (trabucco2021conservative, COMs) is
@@
where $(\cdot)_{\perp}$ is the stop-gradient operator and $\mu_{\theta_{\perp}}({\mathbf{x}})$ is the distribution obtained with a few gradient ascent steps on ${\mathbf{x}}$ initialized from $\mathcal{D}$. This distribution depends on the value of $\theta$ but is not differentiated through while computing the loss’s gradient, and thus we denote it by $\theta_{\perp}$. Unfortunately, in addition to the extra computational cost that the inner-loop gradient ascent induces, COMs’s regularizer limits the amount of improvement that it allows its designs to make.
@@
This is particularly frustrating since recent work on functional graphical models (grudzien2024functional, FGM) delivered a premise of large improvements in the case when the black-box function’s graph can be discovered, as we explain in the next subsection.
@@ 2.2 Functional Graphical Models
An FGM of a high-dimensional function $f({\mathbf{x}})$ is a graph over individual components of ${\mathbf{x}}$ that separates ${\textnormal{x}}_{i},{\textnormal{x}}_{j}\in{\mathbf{x}}$ if their contributions to $f({\mathbf{x}})$ are independent of each other. Knowing such a structure of $f$ one can eliminate interactions between independent variables from a model that approximates it, and thus prevent an MBO algorithm from exploiting them. We summarize basic properties of FGMs below. In what follows, we denote $\mathcal{X}_{-i}$ as the design (input) space without the $i^{th}$ subspace, and ${\mathbf{x}}_{-i}$ as its element  [^1] For example, if $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\mathcal{X}_{3}$ and ${\mathbf{x}}=({\textnormal{x}}_{1},{\textnormal{x}}_{2},{\textnormal{x}}_{3})$, then $\mathcal{X}_{-2}=\mathcal{X}_{1}\times\mathcal{X}_{3}$, and ${\mathbf{x}}_{-2}=({\textnormal{x}}_{1},{\textnormal{x}}_{3})$.. We also write $[K]$ to denote the set $\{1,\dots,K\}$.
@@
Let ${\mathbf{x}}=({\textnormal{x}}_{v}\ |\ v\in\mathcal{V})$ be a joint variable with index set $\mathcal{V}$, and $f({\mathbf{x}})$ be a real-valued function. An FGM $\mathcal{G}=(\mathcal{V},\mathcal{E})$ of $f({\mathbf{x}})$ is a graph where the edge set $\mathcal{E}\subset\mathcal{X}^{2}$ is such that,
$\exists\ f_{-i}:\mathcal{X}_{-i}\rightarrow\mathbb{R}\ \text{and}\ \ f_{-j}:\mathcal{X}_{-j}\rightarrow\mathbb{R},\ \text{with}\ f({\mathbf{x}})=f_{-i}({\mathbf{x}}_{-i})+f_{-j}({\mathbf{x}}_{-j}),\text{implies}\ (i,j)\notin\mathcal{E}.$
See Figure 2 for illustration.
The basic result about FGMs is that they allow for decomposition of the target function into sub-functions with smaller, partially-overlapping inputs, from the FGM’s set of maximal cliques $\mathcal{C}$,
@@
Theorem 1 (grudzien2024functional).
Let $f({\mathbf{x}})$ be a real-valued function, $\mathcal{C}$ be the set of maximal cliques of its FGM, and $\Pi$ be a policy class. Let $\text{C}_{\text{stat}}$ and $\text{C}_{\text{cpx}}$ be constants that depend on the probability distribution of ${\mathbf{x}}$ and function approximator class’s complexity, respectively, defined in Appendix A. Then, the regret of MBO with the FGM information is given by,
$\eta(\pi^{\star})-\eta(\hat{\pi}_{\text{FGM}})\leq\text{C}_{\text{stat}}\text{C}_{\text{cpx}}\max\limits_{\pi\in\Pi,{\mathbf{x}}\in\mathcal{X},C\in\mathcal{C}}\frac{\pi({\mathbf{x}}_{C})}{p_{C}({\mathbf{x}}_{C})}$.
The implication of this theorem is that the FGM-equipped function approximator does not require the dataset to cover the entire design space. Rather, it only requires that the individual cliques of the space be covered, which is a much milder requirement, especially when the cliques are small. In the next section, we show how these results can be combined with a transformer, mitigate the distribution shift problem, and enable efficient MBO.
@@ 3 Cliqueformer
This section introduces a neural network model designed to solve MBO problems through standard end-to-end training on offline datasets. We present a new theoretical result, outline the key desiderata for such a model, and propose an architecture—Cliqueformer—that addresses these requirements.
3.1 Non-Uniqueness of Structure Discovery
Figure 3: Illustration of construction in the proof of Theorem 2 for $d=2$. Red axes represent ${\mathbf{z}}$ and blue axes represent ${\mathbf{v}}$. When considered a function of ${\mathbf{z}}$, the contour curves (straight lines) of $f$ are a function of both ${\textnormal{z}}_{1}$ and ${\textnormal{z}}_{2}$, but as a function of ${\mathbf{v}}$, they only depend on ${\textnormal{v}}_{1}$. The density of the Gaussian distribution (green curves) are identical circles for both variables.
The regret bound from Theorem 1 applies to methods that use the target function’s FGM in their function approximation. It implies that such methods can solve even very high-dimensional problems if their underlying FGMs have low-dimensional cliques or, simply speaking, are sparse. Since, in general, no assumptions about the input can be made, this motivates learning a representation of the input for which one can make distributional assumptions and infer the FGM with statistical tests. Following this reasoning, grudzien2024functional offer a heuristic technique for discovering an FGM over learned, latent, normally-distributed variables. However, as we formalize with the following theorem, even such attempts are futile in dealing with black-box functions.
@@ 3.1 Non-Uniqueness of Structure Discovery
Let $d\geq 2$ be an integer and ${\mathbf{x}}\in\mathbb{R}^{d}$ be a random variable with positive density in $\mathbb{R}^{d}$. There exists a function $f({\mathbf{x}})$ and two different reparameterizations, ${\mathbf{z}}=z({\mathbf{x}})$ and ${\mathbf{v}}=v({\mathbf{x}})$, of ${\mathbf{x}}$, that both follow a standard-normal distribution, but the FGM of $f$ with respect to ${\mathbf{z}}$ is a complete graph (has all possible edges), and with respect to ${\mathbf{v}}$ it is an empty graph (has no edges).
Proof Sketch.
The proof is a construction. We first map ${\mathbf{x}}$ to a standard-normal variable ${\mathbf{z}}$ with tools from high-dimensional statistics. We then introduce scalar variable y that is a function of ${\mathbf{z}}$ and has a complete FGM. We then show that we can rotate ${\mathbf{z}}$ onto another normal variable ${\mathbf{v}}$ with respect to which y is a function with an empty FGM (see Figure 3 for illustration). We complete the proof by showing how to express y as a function of ${\mathbf{x}}$. The full proof can be found in Appendix A.
The theorem implies that FGM is not a fixed attribute of a function that can be estimated from the data, but instead should be viewed as a property of the input’s reparameterization. Furthermore, different reparameterizations feature different FGMs with varied levels of decomposability, some of which may not significantly simplify the target function. This motivates a reverse approach that begins from defining a desired FGM and learning representations of the input that align with the graph. In the next subsection, we introduce Cliqueformer, where the FGM is specified as a hyperparameter of the model and a representation of the data that follows its structure is learned.

### v2

@@ Abstract
Large neural networks excel at prediction tasks, but their application to design problems, such as protein engineering or materials discovery, requires solving offline model-based optimization (MBO) problems. While predictive models may not directly translate to effective design, recent MBO algorithms incorporate reinforcement learning and generative modeling approaches. Meanwhile, theoretical work suggests that exploiting the target function’s structure can enhance MBO performance. We present Cliqueformer, a transformer-based architecture that learns the black-box function’s structure through functional graphical models (FGM), addressing distribution shift without relying on explicit conservative approaches. Across various domains, including chemical and genetic design tasks, Cliqueformer demonstrates superior performance compared to existing methods.
@@ 1 Introduction
Most of the common use cases of deep learning (DL) so far have taken the form of prediction tasks (Hochreiter & Schmidhuber, 1997; He et al., 2016; Krizhevsky et al., 2017; Vaswani et al., 2017). However, in many applications, such as protein synthesis or chip design, we might want to use powerful models to instead solve optimization problems. Clearly, accurate predictions of a target score of an object could be used to find a design of the object that maximizes that score. Such a methodology is particularly useful in engineering problems in which evaluating solution candidates comes with big risk. For example, synthesizing a proposed protein requires a series of wet lab experiments and induces extra cost and human effort (Gómez-Bombarelli et al., 2018; Brookes et al., 2019). Thus, to enable proposing de-novo generation of strong solution candidates, researchers have drawn their attention to offline black-box optimization (BBO), often referred to as model-based optimization (MBO). In this paradigm, first, a surrogate model of the score is learned from offline data. Next, a collection of designs is trained to maximize the surrogate, and then proposed as candidates for maximizers of the target score (Gómez-Bombarelli et al., 2018; Kumar et al., 2021).
@@
Figure 1: The first building block of the Lat. RBF tasks are 3D radial-basis functions (left). These functions are applied to triplets arranged in a chain of triangles FGM (center) and linearly mixed. Then, observable designs are produced with non-linear transformations of the chain and, together with their values, form a dataset. We show the score (right) of our structure-learning Cliqueformer and structure-oblivious COMs (Trabucco et al., 2021), against the dimension of Lat. RBF functions, modulated only by varying the number of triangles. Cliqueformer, unlike COMs, sustains strong performance across all dimensions. More results in Section 5.
In contrast to previous works, in our paper, we develop a scalable model that tackles MBO by learning the structure of the black-box function through the formalism of functional graphical models. Our architecture aims to solve MBO by 1) decomposing the predictions over the cliques of the function’s FGM, and 2) enforcing the cliques’ marginals to have wide coverage with our novel form of the variational bottleneck (Kingma & Welling, 2013; Alemi et al., 2016). However, building upon our Theorem 2, we do not follow Grudzien et al. (2024) during the FGM discovery step, and instead subsume it in the learning algorithm. To enable scaling to high-dimensional problems and large datasets, we employ the transformer backbone (Vaswani et al., 2017). Empirically, we show that our model, Cliqueformer, inherits the scalability guarantees of MBO with FGM (see Figure 1). We further demonstrate its superiority to baselines in a suite of tasks with latent radial-basis functions (Grudzien et al., 2024) and real-world chemical (Hamidieh, 2018) and DNA design tasks (Trabucco et al., 2022; Uehara et al., 2024a).
@@ 2 Preliminaries
In this section, we provide the necessary background on offline model-based optimization. Additionally, we cover the basics of functional graphical models on top of which we build Cliqueformer.
2.1 Offline Model-based Optimization
We consider an offline model-based optimization (MBO for short) problem, where we are given a dataset $\mathcal{D}=\{{\mathbf{x}}^{i},{\textnormal{y}}^{i}\}_{i=1}^{N}$ of examples ${\mathbf{x}}\in\mathcal{X}$, following distribution $p({\mathbf{x}})$, and their values ${\textnormal{y}}=f({\mathbf{x}})\in\mathbb{R}$ under an unkown (black-box) function $f:\mathcal{X}\rightarrow\mathbb{R}$. Our goal is to optimize this function offline. That is, to find its maximizer
@@ 2.1 Offline Model-based Optimization
by only using information provided in $\mathcal{D}$ (Kumar & Levine, 2020). Sometimes, a more general objective in terms of a policy over $\pi({\mathbf{x}})$ is also used, $\eta({\mathbf{x}})=\mathbb{E}_{{\mathbf{x}}\sim\pi}[f({\mathbf{x}})]$. In either form, unlike in Bayesian optimization, we cannot make additional queries to the black-box function (Brochu et al., 2010; Kumar & Levine, 2020). This formulation represents settings in which obtaining such queries is prohibitively costly, such as tests of new chemical molecules or of new hardware architectures (Kim et al., 2016; Kumar et al., 2021; Yang et al., 2024).
@@
where $\text{Reg}(\theta)$ is an optional regularizer. Classical methods choose $\text{Reg}(\theta)$ to be identically zero, while conservative methos use the regularizer to bring the values of examples out of $\mathcal{D}$ down. For example, the regularizer of Conservative Objective Model’s (Trabucco et al., 2021, COMs) is
@@
where $(\cdot)_{\perp}$ is the stop-gradient operator and $\mu_{\theta_{\perp}}({\mathbf{x}})$ is the distribution obtained with a few gradient ascent steps on ${\mathbf{x}}$ initialized from $\mathcal{D}$. This distribution depends on the value of $\theta$ but is not differentiated through while computing the loss’s gradient, and thus we denote it by $\theta_{\perp}$.
@@
Unfortunately, in addition to the extra computational cost that the inner-loop gradient ascent induces, COMs’s regularizer limits the amount of improvement that it allows its designs to make. This is particularly frustrating since recent work on functional graphical models (Grudzien et al., 2024, FGM) delivered a premise of large improvements in the case when the black-box function’s graph can be discovered, as we explain in the next subsection.
@@ 2.2 Functional Graphical Models
An FGM of a high-dimensional function $f({\mathbf{x}})$ is a graph over components of ${\mathbf{x}}$ that separates ${\textnormal{x}}_{i},{\textnormal{x}}_{j}\in{\mathbf{x}}$ if their contributions to $f({\mathbf{x}})$ are independent of each other. Knowing such a structure of $f$ one can eliminate interactions between independent variables from a model that approximates it, and thus prevent an MBO algorithm from adversarially exploiting them. We summarize basic properties of FGMs below. In what follows, we denote $\mathcal{X}_{-i}$ as the design (input) space without the $i^{th}$ subspace, and ${\mathbf{x}}_{-i}$ as its element  [^1] For example, if $\mathcal{X}=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\mathcal{X}_{3}$ and ${\mathbf{x}}=({\textnormal{x}}_{1},{\textnormal{x}}_{2},{\textnormal{x}}_{3})$, then $\mathcal{X}_{-2}=\mathcal{X}_{1}\times\mathcal{X}_{3}$, and ${\mathbf{x}}_{-2}=({\textnormal{x}}_{1},{\textnormal{x}}_{3})$.. We also write $[K]$ to denote the set $\{1,\dots,K\}$.
@@
Let ${\mathbf{x}}=({\textnormal{x}}_{v}\ |\ v\in\mathcal{V})$ be a joint variable with index set $\mathcal{V}$, and $f({\mathbf{x}})$ be a real-valued function. An FGM $\mathcal{G}=(\mathcal{V},\mathcal{E})$ of $f({\mathbf{x}})$ is a graph where the edge set $\mathcal{E}\subset\mathcal{X}^{2}$ is such that, $\exists\ f_{-i}:\mathcal{X}_{-i}\rightarrow\mathbb{R}\ \text{and}\ \ f_{-j}:\mathcal{X}_{-j}\rightarrow\mathbb{R},\ \text{with}\ f({\mathbf{x}})=f_{-i}({\mathbf{x}}_{-i})+f_{-j}({\mathbf{x}}_{-j}),\text{implies}\ (i,j)\notin\mathcal{E}.$ See Figure 2 for illustration.
Crucially, FGMs enable decomposing the target function into sub-functions with smaller inputs defined by the set of maximal cliques $\mathcal{C}$ of the FGM’s graph,
@@
Theorem 1 (Grudzien et al. (2024)).
For a real-valued function $f({\mathbf{x}})$ with FGM with maximal cliques $\mathcal{C}$, and policy class $\Pi$, the regret of MBO with FGM information satisfies $\eta(\pi^{\star})-\eta(\hat{\pi}_{\text{FGM}})\leq\text{C}_{\text{stat}}\text{C}_{\text{cpx}}\max\limits_{\pi\in\Pi,{\mathbf{x}}\in\mathcal{X},C\in\mathcal{C}}\frac{\pi({\mathbf{x}}_{C})}{p_{C}({\mathbf{x}}_{C})}$ where $\text{C}_{\text{stat}}$, $\text{C}_{\text{cpx}}$ are distribution and complexity constants defined in Appendix B.
This theorem implies that a function approximator equipped with FGM does not require the dataset to cover the entire design space. Instead, it suffices for the dataset to span individual cliques of the space, which is a substantially less stringent condition—particularly when the cliques are small. To illustrate this result numerically, we conduct an experiment, shown in the left column of Figure 3. Specifically, we generate high-dimensional Gaussian data ${\mathbf{x}}$ and evaluate $f({\mathbf{x}})$, which is modeled as a mixture of random radial basis functions (RBFs) constrained by an FGM of triangles, similar to the setup in Figure 1. We then train two models, $f_{\theta}({\mathbf{x}})$, each with the same parameter count: one adhering to the FGM decomposition and the other not. These models are used both to approximate the target function and to optimize ${\mathbf{x}}$. The results demonstrate that the FGM-equipped model significantly outperforms its counterpart, both in terms of the quality of fit and, more critically, in the quality of the optimized designs.
(a) MSE loss w/ known FGM
(c) MSE loss w/o known FGM
(b) f(x) value w/o known FGM
(d) f(x) value w/o known FGM
Figure 3: Left column. FGM known. The model without the FGM information (red) fits the data poorly (Fig. (3(a))) and leads to poor designs (Fig. (3(b))). The model with FGM information (blue) achieves a good fit and leads to designs largely superior to the data. Righ column. FGM unknown. The model with an FGM decomposition (blue) achieves a slightly better fit than an oblivious model (red) but leads to significantly better designs.
In the next section, we show how these results can be combined with a transformer, mitigate the distribution shift problem, and enable efficient MBO.
@@ 3 Cliqueformer
This section introduces a neural network model to solve MBO problems through standard end-to-end training on offline datasets. We present a new theoretical result, outline the key desiderata for such a model, and propose an architecture—Cliqueformer—that addresses these requirements.
3.1 Structure Discovery
The regret bound from Theorem 1 applies to methods that use the target function’s FGM in their function approximation. It implies that such methods can solve even very high-dimensional problems if their underlying FGMs have low-dimensional cliques or, simply speaking, are sparse. Since, in general, no assumptions about the input can be made, this motivates learning a representation of the input for which one can make distributional assumptions and infer the FGM with statistical tests. Following this reasoning, Grudzien et al. (2024) offer a heuristic technique for discovering an FGM over learned, latent, normally-distributed variables. However, as we formalize with the following theorem, which we prove in Appendix B, even such attempts are futile in dealing with black-box functions.
@@ 3.1 Structure Discovery
For any $d\geq 2$ and random variable ${\mathbf{x}}\in\mathbb{R}^{d}$ with positive probability density, there exists a function $f({\mathbf{x}})$ and two standard-normal reparameterizations ${\mathbf{z}}=z({\mathbf{x}})$, ${\mathbf{v}}=v({\mathbf{x}})$ such that the FGM of $f$ is complete with respect to ${\mathbf{z}}$ but empty with respect to ${\mathbf{v}}$.
The theorem (see Figure 4 for intuition behind the proof) implies that FGM is not a fixed attribute of a function that can be estimated from the data, but instead should be viewed as a property of the input’s reparameterization. Furthermore, different reparameterizations feature different FGMs with varied levels of decomposability, some of which may not significantly simplify the target function.
Figure 4: Illustration of construction in the proof of Theorem 2 in 2D. Red axes (${\mathbf{z}}$) show contour lines of $f$ depending on both coordinates ${\textnormal{z}}_{1}$ and ${\textnormal{z}}_{2}$. Blue axes (${\mathbf{v}}$) show the same contours depending only on ${\textnormal{v}}_{1}$ after rotation. The Gaussian distribution (green circles) maintains circular shape in both coordinate systems, demonstrating invariance under rotation.
This motivates a reverse approach that begins from defining a desired FGM and learning representations of the input that align with the graph. To examine if FGM decomposition implemented this way can still bring benefits in MBO, we conduct another experiment in the right column of Figure 3. This time, we conceal the FGM information from both models, while implementing one of them with another decomposition, in a latent space that is computed with an attention layer (for fairness, we gave the FGM-oblivious model an attention layer too). Despite not bringing major improvements to the quality of fit, the decomposable model does lead to better designs, providing empirical support for the considered approach. Consequently, in the next subsection, we introduce Cliqueformer, where the FGM is specified as a hyperparameter of the model and a representation of the data that follows its structure is learned.